At the local drive-in theater, the manager sold 200 tickets. Adult tickets cost $9 each, while children's tickets cost $5 each. If the theater sold $1420 worth of tickets, how many adult tickets and how many children's tickets were sold?



Found 3 solutions by mananth, Theo, greenestamps:
Answer by mananth(16946)

At the local drive-in theater, the manager sold 200 tickets. Adult tickets cost $9 each, while children's tickets cost $5 each. If the theater sold $1420 worth of tickets, how many adult tickets and how many children's tickets were sold?
Let number of adult tickets sold be x
and the number of child tickets be y
the manager sold 200 tickets.
x+y=200---------------------------------1
Adult tickets cost $9 each, while children's tickets cost $5 each. If the theater sold $1420 worth of tickets
Revenue from sales
9x+5y =1420------------------------------2
Multiply (1) by 5 and subtract from (2)
9x+5y =1420
5x+5y=1000
4x = 420
x = 420/4
x= 105 the number of adult tickets sold
you continue




Answer by Theo(13342)

a = number of adult tickets
c = number of child tickets.
your equations are:
a + c = 200
9a + 5c = 1420
multiply boh sides of the first equation by 9 and leave the second equation as is to get:
9a + 9c = 1800
9a + 5c = 1420
subtract the second equation from the first to get:
4c = 380
solve for c to get:
c = 380/4 = 95
a = 200 - 95 = 105
you have 105 adult tickets and 95 child tickets.
total cost is 9 * 105 + 5 * 95 = 1420.
numbers look good.
your solution is 105 adult and 95 child tickets were sold.

Answer by greenestamps(13200)



The formal algebraic solutions provided by the other tutors are fine.

But you can get good mental exercise, and good problem-solving practice, by solving the problem using logical reasoning and simple mental arithmetic.

If all 200 tickets were children's tickets, the total cost would be 200($5) = $1000.
The actual total cost was $1420-$1000 = $420 more than that.
Each adult ticket costs $4 more than each children's ticket.
The number of adult tickets sold was $420/$4 = 105.

ANSWER: 105 adult tickets and 200-105 = 95 children's tickets.